OpenStudy (anonymous):

Line L1 passes through point A, and is PARALLEL to line L2 that passes through P and Q. A(0,2); P(1,4) and Q(3,4) Slope of L2 Blank 1 Slope of L1 Blank 2 y-intercept of L1 (as a point B) Blank 3 L1 in standard form Blank 4

OpenStudy (anonymous):

First find slope for L2 and L2 and L1 are PARALLEL so : m1=m2 m2&m2=y2-y1/x2-x1=>4-4/3-1=0/2=0 ______________________________________________________ y-intercept of L1 (as a point B) : Which one is point B ?!@aleepanda